OpenStudy (anonymous):

2x^2-5x=-10 It ask for the solution set is Use the quadratic formula to solve the equation

OpenStudy (anonymous):

I provided the formula in the other problem you had. 2x^2-5x+10=0 a=2 b=-5 c=10 (5+-sqrt(25-80))/2. You can't find the square root of a negative number without using complex numbers (which, I guess, is way beyond your level), so the right answer is "no answer".

OpenStudy (anonymous):

it is not giving me that option, it says there is an answer?

OpenStudy (dumbcow):

try \[\frac{5\pm \sqrt{55}i}{4}\] otherwise make sure you copied the question correctly

OpenStudy (anonymous):

ok I checked it the problem is copied correctly. It ask what the solution set is \[\left\{ ? \right\}\] Then under that it says, type an exact answer, using radicals as needed. Express complex numbers in terms of i. Use integers or fractions for any numbers in the expression. Use comma to seperate answers

OpenStudy (dumbcow):

ok above answer is correct

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Okay, my bad then. I figured you didn't have to provide a complex answer.